- Gemini Home
- Telescopes and Sites
- Science Visitors at Gemini
- Observing With Gemini
- Retired Instruments
- Visiting Instrument Policy
- Visiting Instrument Telescope Interfaces
- DSSI Speckle Camera
- TEXES (North)
- Integration Time Calculators
- Magnitudes and Fluxes
- Near-IR Resources
- Mid-IR Resources
- Observing Condition Constraints
- Performance Monitoring
- SV/Demo Science
- Future Instrumentation & Current Development
- Queue and Schedules
- Data and Results
- Gemini Research Staff
Change page style:
NIRI Update: mechanisms examined and fixed
7 October 2002
NIRI has been opened and the faulty mechanisms examined.
- The stepper motor on the pupil wheel failed. The stepper motor was replaced with a spare, and the mechanism thoroughly checked. No other damage was found, and the wheel is working again.
- The OIWFS gimbal mirror had in fact run through a limit switch without stopping. The switch was damaged, as were some of the fragile linkages in the gear train. The damaged parts were replaced with spares, and the gears carefully inspected. The mechanical parts of the mechanism are working once again.
- The electronics associated with the OIWFS gimbal mirror sensors and limit switches were examined. One cable connector inside the external electronics cabinet was destroyed, probably by a short between a power supply and the ground shield on the cable. The cable and connector were replaced and the electronics tested. The short was almost certainly the cause of the sensor failure that led to the OIWFS gimbal mirror running through its limit switch.
NIRI will be cooled again in the lab during the coming week. We will then enter a period of mechanism testing and instrument engineering in support of Altair commissioning in mid-November. After the lab testing, the NIRI mechanisms will be left in position for the Altair commissioning.
If all goes well with the lab testing, instrument engineering, and Altair commissioning, we may be able to resume queue observations sometime in December. No additional Phase II submisions will therefore be needed until the first of December. Please watch these pages for more information on the results of the testing and plans for resuming queue observations.
Last update Oct. 7, 2002; Joe Jensen